wonder if i am at the right place if not please direct me tell me is this correct please (-12)+5-(-7)=-12+5+7=1-12+12=0 this is what my teacher say i cant make out how it goes thank you.
Answer by bonster(299)

(-12)+5-(-7)=-12+5+7=1-12+12=0

(-12)=-12
+5=+5
-(-7)=+7

so:
(-12)+5-(-7)
=-12+5+7
=-12+12
=0